OpenStudy (anonymous):

find the gradients of each of these lines 1) y=4-3x 2)2y-2x=4

OpenStudy (saifoo.khan):

Slope is the coefficent of x.

OpenStudy (saifoo.khan):

For 1) -3 2) 1

OpenStudy (anonymous):

saifooo can you please show workings for number 2

OpenStudy (anonymous):

the gradient of a line is the slope 2y-2x=4 2y=2x+4 add 2x both sides y=(2x+4)/2 divide by 2 both sides y=x+2 simplify each term (2x/2=x and 4/2=2) the slope (gradient) is 1

OpenStudy (anonymous):

Thank you